# Physics-Driven Construction of Compact Primitive Gaussian Density Fitting Basis Sets

### v1

@@ Abstract
We present model-assisted density fitting (MADF) basis set generator, an algorithm for generating primitive atomic Gaussian density fitting (DF) basis sets (DFBSs) from a contracted Gaussian orbital basis set (OBS). The MADF algorithm produces DFBSs suitable for accurate robust DF approximation of 2-particle interactions in mean-field and correlated electronic structure. The algorithm is designed to (a) saturate the OBS product space by a large regularized set of primitive solid-harmonic Gaussian shells with nonuniform distribution of exponents followed by (b) pruning of the shells according to their contributions to the 2-body energy of a correlated atomic ensemble. Building the DFBS generator model almost exclusively on mathematical and physical principles allows one to limit the number of parameters that control the density fitting error to four, with a single set of parameters sufficient for computations with all basis cardinal numbers, with and without correlation of core electrons, with and without scalar and spin-dependent relativistic effects, spanning almost all of the Periodic Table. Performance assessment included basis sets up to quadruple-zeta quality from several major basis set families, using molecules composed of main-group, d-block, and f-block elements. The resulting DF errors in Hartree-Fock and second-order MP2 energies (with relativistic all-electron treatments, when appropriate) were on the order of $20$ and $10$ $\mu E_{h}$ per electron, respectively.
@@ 1 Introduction
Most DBFS generators are designed only for DF in the context of evaluating the Coulomb potential of a charge density. These include the AutoABS method by Yang et al.57, long-range corrected auxiliary basis sets by Hellmann et al.58, even tempered auxiliary basis sets with shared exponents by Daz-Tinoco et al.59 and a similar approach has also recently been reported for relativistic OBSs60.
@@
The AutoAux method, available in the ORCA software package64 and in the command-line interface of Basis Set Exchange (BSE)41, produces a primitive even-tempered DFBS that can be applied universally (in mean-field and correlated simulations). The AutoAux generator uses even-tempered sequences of Gaussian primitive AOs to span the OBS-deduced range of exponents for each angular momentum block. The exponent ranges and even-tempered ratios vary with the angular momenta ($L$) and the atomic number $Z$ (these even-tempering parameters are prescribed for up to $L=7$; the BSE implementation of AutoAux uses the same exponent ratio for $L\geq 7$). The generator also limits the highest angular momentum of DFBS AOs below what is needed for exact atomic DF, in a manner that depends on the highest angular momentum of the occupied orbitals. AutoAux as a result is quite complex, defined with more than a dozen model parameters in total.
@@
PySCF’s simple built-in generator65 produces even-tempered primitive DFBS and is very similar in spirit to AutoAux. It has fewer model parameters (e.g., even-tempering ratio is not graded with angular momenta) and claims to offer similar accuracy to AutoAux, although no extensive comparison exists.
@@
Lehtola presented a DFBS generator that produces nearly exact atomic DFBS composed of primitive solid-harmonic Gaussian (SHG) AOs.62 First, the OBS AOs on a given atom are used to generate a base pool of primitive SHGs by approximating each angular momentum channel of solid-harmonic OBS AO products by a single solid-harmonic AO. The resulting pool is then regularized by pivoted Cholesky decomposition (pCD) of the two-center two-electron Coulomb integral matrix, with the accuracy of the resulting primitive DFBS controlled robustly by a single threshold. Although the resulting DFBSs are nearly exact, they are extremely large, lead to high condition numbers in molecules, and contain primitive Gaussians with twice the angular momentum of the highest angular momentum ($L_{\text{OBS}}$) of functions in the parent OBS. Lehtola addressed the latter issue in his recent work66 by extending the algorithm to prune high-angular momentum functions out of the DFBS with a tunable parameter. The algorithm then uses Kállay’s67 contraction method with singular value decomposition (SVD) to contract the DFBS, further reducing its size.
An attractive alternative to density fitting that avoids these issues and provides some of its advantages is the (pivoted) Cholesky decomposition (CD).68 Any positive-definite 2-particle interaction represented in an arbitrary OBS can be decomposed with accuracy controlled by a single threshold. CD has been used for electronic structure simulations in both relativistic69, 70, 71 and nonrelativistic72, 73, 74 contexts. Although the accuracy of CD can be robustly pushed beyond the reach of conventional density fitting, the cost of CD is substantial. Computing analytic derivatives of the Cholesky vectors is also fundamentally challenging. The 1-center CD (1C-CD),75 later renamed atomic CD (aCD),76 works around the problem of analytic derivatives of CD by marrying CD with density fitting. Namely, global CD is avoided and CD is only used to construct a contracted DFBS for each atom suitable for accurate representation of a given operator in a given OBS. Thus the aCD approach can be viewed as a black-box 1-parameter DFBS generator. The aCD approach also produces contracted DFBSs which have mixed angular momentum functions and mixed representations (cartesian or spherical). A solution to this issue was proposed by Hellmann et al.77, combining aCD and Lehtola’s method62, however, it suffers from the same issues as Lehtola’s original method.62 Several challenges in designing efficient CD algorithms in the context of quantum chemistry have also been mentioned.78
@@
To navigate the severe gaps in the coverage by manually-constructed DFBS of existing work and the significant shortcomings of the existing DFBS generators we attempted to design a new DFBS generator that produces (a) primitive Gaussian DFBSs, (b) of comparable size and accuracy as the manually-optimized DFBS, (c) usable in the context of mean-field and correlated simulation, and (d) has as few adjustable parameters as possible to ensure maximum universality. By combining the existing ideas for spanning the product space of OBS AOs with new ideas for regularizing and pruning the exact DFBS using atomic correlated ensemble RDMs we arrived at a DFBS generator dubbed MADF. The purpose of this manuscript is to describe it and assess its performance. Following a quick recap of density fitting in Section 2.1 we describe the 2 key ingredients of the MADF generator: the regularization of the complete set of candidate primitives (described in Section 2.2) and subsequent pruning by a 2-body energy estimator model (described in Section 2.3). Section 3 describes the technical and details of the parameter training regiment. In Section 4 we document the performance of the DFBSs generated by our method against the manually-optimized DFBSs and the DFBSs generated by the AutoAux algorithm.
@@ 2.1 Density Fitting
$|\mu\nu)\approx\sum_{X}C_{\mu\nu}^{X}|X)\equiv|\widetilde{\mu\nu}).$ (1)
The fitting coefficients $C_{\mu\nu}^{X}$ are obtained by minimizing a norm of the density error $|\delta_{\mu\nu})\equiv|\mu\nu)-|\widetilde{\mu\nu})$. To minimize the error in the diagonal elements of the matrix representation of a positive operator $\hat{O}$, namely in $(\mu\nu|\hat{O}|\mu\nu)$, the optimal choice of the error norm is the expectation value of operator $\hat{O}$ itself:
$||f||_{\hat{O}}\equiv(f|\hat{O}|f),$ (2)
This is known as robust density fitting.79, 80 Least-squares minimization of $||\delta_{\mu\nu}||_{\hat{O}}$ is equivalent to solving the following equation
@@
DFBS generators usually57, 58, 59, 60, 61, 62, 63, 64 e.g., AutoAux, aCD work in a single-atom context, i.e., DFBS is constructed using information about products on OBS AOs on a single atom. It is reasonable to question whether this is sufficient. Assuming a system composed of a single atom type, global density fitting in a system of $n$ atoms will be able to fit every such 1-center product of OBS AOs using $n$ times as many DFBS AOs as in a single atom. Indeed, for long-range $\hat{O}$ the fitting of even strongly-localized AO products generally involves DFBS AOs from far away.81 Thus, 1-center products in a molecule can borrow DFBS AOs from other centers to make the fitting more accurate than in a single atom. This suggests that a single atom poses a more challenging DF setting than a molecule, which supports the use of a single-atom setting for the generation of DFBS.
@@ 2.2 Products of primitive Gaussians AOs
The density fitting becomes exact if the span of DFBS matches the span of the OBS AO products. Because DFBS AOs are atom-centered but the products of AOs in a molecule are generally centered between atoms, exact DF is only possible for a single atom. We then confine ourselves to a single atom. Furthermore, we will limit ourselves to the case of OBS and DFBS composed of solid-harmonic Gaussian (SHG) AOs. Unnormalized primitive SHG with exponent $\alpha$ is defined as:82, 83
@@
Whereas product of any number of concentric primitive Cartesian Gaussians is a single Cartesian Gaussian (modulo normalization), product of concentric SHGs $\chi_{\alpha_{\mu},l_{\mu},m_{\mu}}$ and $\chi_{\alpha_{\nu},l_{\nu},m_{\nu}}$ is a linear combination of $r^{l_{\mu}+l_{\nu}-L}\chi_{\alpha_{\mu}+\alpha_{\nu},L,m_{\mu}+m_{\nu}}$ for $\max(|l_{\mu}-l_{\nu}|,m_{\mu}+m_{\nu})\leq L\leq l_{\mu}+l_{\nu}$. Clearly, the contribution to the product from the Clebsch-Gordan channel $L<l_{\mu}+l_{\nu}$ is an SHG multiplied by $r^{l_{\mu}+l_{\nu}-L}$. Lehtola proposed to account for this effect by approximating each such contribution by a single SHG of angular momentum $L$ with the following effective exponent (see Appendix II of Ref. 62):
@@
Let $\mathcal{C}$ denote the “complete” set of all unique solid harmonic primitive AOs generated thereby from all products of primitive OBS AOs on a given atom. The set $\mathcal{C}$ can be used as DFBS for nearly exact density fitting in an atom (a) for any contraction of the generating set of primitive OBS AOs and (b) for any operator. However, even in a single atom such DFBS is highly linearly dependent and overcomplete, 84, 62 therefore, it is unusable with finite precision arithmetic. Moreover, such a basis would be extremely large, defeating the purpose of the DF approximation. Hence, a compact DFBS is needed to represent the product space of AO functions. Since the set $\mathcal{C}$ is overcomplete for use as a DFBS for an atom, the unimportant and linearly dependent functions must be trimmed out of this set. Lehtola proposed to do such pruning by pCD.62 However, even after pCD the resulting DFBSs are still large, as illustrated in Figure 1. While pCD reduces the density of the complete set, it largely leaves its range of exponents intact (at least for lower $L$), and it still keeps many pairs of adjacent exponents with small ratios ($\ll 2$). The problem is especially pronounced for the high angular momentum ($L=6$).
@@
Figure 1: Exponents of DFBS generated (or manually-optimized) for Kr atom with representative nonrelativistic and relativistic OBS using various methods scattered on a logarithmic number line for different angular momenta i.e., $L=\{0,1,2,6\}$. The number of exponents is shown on the right of each scatter plot. Figure 1(a) shows distribution of DFBS exponents for nonrelativistic OBS def2-QZVPP and Figure 1(b) for relativistic OBS x2c-QZVPPall-2c. The plot depict Lehtola’s complete set (shown in purple), Lehtola’s pCD-regularized set (shown in red) with pCD threshold $10^{-7}$ as recommended in Ref.62, even-tempered exponents produced by AutoAux (shown in green) and set of exponents produced by the MADF algorithm (shown in orange). Distribution of exponents of manually-optimized DFBS def2-QZVPP-RIFIT is also shown (in blue) in Figure 1(a).
@@
In this work, we regularize the complete candidate exponent pool using a simple adaptive algorithm designed to produce non-even-tempered sets that deviate from the input set as little as possible and with exponent ratios as least as large as the target ratio threshold. Given a primitive SHG set $\mathcal{C}$ and a target exponent ratio threshold $\zeta$, each subset $\mathcal{C}_{L}$ containing primitives of angular momenta $L$ is regularized by repeatedly replacing the pair of primitive SHG shells with the smallest exponent ratio $\alpha_{1}/\alpha_{2}$ by a single primitive SHG shell with the geometric mean exponent $\alpha=\sqrt{\alpha_{1}\alpha_{2}}$. Regularization stops if no pairs of primitives have an exponent ratio less than $\zeta$. Clearly, regularization is most easily implemented if $\mathcal{C}_{L}$ is first sorted by the exponent in descending order. The algorithm is described in detail in Algorithm 1. The recommended value of $\zeta$ will be determined heuristically in Section 4.1.
@@
2:     $\mathcal{C}\leftarrow\textsc{CompleteDFBS}(\mathcal{P})$triangleright Complete set of DF SHG shells with exponents generated by Eq. 8
3:     $L_{\text{max}}\leftarrow\textsc{MaxAM}(\mathcal{c})$triangleright Maximum angular momentum of shells in $\mathcal{c}$
4:     ${}\mathcal{R}\leftarrow\{\}$triangleright Initialize the regularized complete DFBS
@@
6:         $\mathcal{c}\leftarrow\textsc{SelectByAM}(\mathcal{C},L)$triangleright Shells in $\mathcal{C}$ with angular momentum $L$
7:         if Size($\mathcal{c}$)$\geq 2$ then
8:              $\zeta_{\text{min}}\leftarrow\infty$triangleright smallest exponent ratio
@@
11:                  $\mathcal{c}\leftarrow\textsc{SortByExponent}(\mathcal{c})$triangleright Sort in order of decreasing exponents
12:                  for $i$ in $20\dots\textsc{Size}(\mathcal{c})-$ 2 do
13:                       $\zeta_{i}\leftarrow\textsc{Exponent}(\mathcal{c}\left[i\right])/\textsc{Exponent}(\mathcal{c}\left[i+1\right])$
14:                       if $\zeta_{i}<\zeta_{\text{min}}$ then
15:                           $\zeta_{\text{min}}\leftarrow\zeta_{i}$
16:                           $I_{\text{min}}\leftarrow i$triangleright Tracks shell pair with smallest exponent ratio
17:                       else
18:                           continue
19:                       end if
20:                  end for
21:                  $\alpha_{I}\leftarrow\textsc{Exponent}(\mathcal{c}\left[I_{\text{min}}\right])$
22:                  $\alpha_{I+1}\leftarrow\textsc{Exponent}(\mathcal{c}\left[I_{\text{min}}+1\right])$
23:                  $c\leftarrow\textsc{SHGShell}(\sqrt{\alpha_{I}\alpha_{I+1}},L)$triangleright Pair of shells with closest exponents fused to single shell
[…]
25:              end while
26:         end if
27:         $\mathcal{R}\leftarrow\textsc{Concat}(\mathcal{R},\mathcal{c})$
28:     end for
29:     return $\mathcal{R}$
30:end function

### v2

@@ Abstract
We present model-assisted density fitting (MADF) basis set generator, an algorithm for generating primitive atomic Gaussian density fitting (DF) basis sets (DFBSs) from a contracted Gaussian orbital basis set (OBS). The MADF algorithm produces DFBSs suitable for accurate robust DF approximation of 2-particle interactions in mean-field and correlated electronic structure. The algorithm is designed to (a) saturate the OBS product space by a large regularized set of primitive solid-harmonic Gaussian shells with nonuniform distribution of exponents followed by (b) pruning of the shells according to their contributions to the 2-body energy of a correlated atomic ensemble. Building the DFBS generator model almost exclusively on mathematical and physical principles allows one to limit the number of parameters that control the density fitting error to three, with a single set of parameters sufficient for computations with all basis cardinal numbers, with and without correlation of core electrons, with and without scalar and spin-dependent relativistic effects, spanning almost all of the Periodic Table. Performance assessment included basis sets up to quadruple-zeta quality from several major basis set families, using molecules composed of main-group, d-block, and f-block elements. The resulting DF errors in Hartree-Fock and second-order MP2 energies (with relativistic all-electron treatments, when appropriate) were on the order of $20$ and 1010 $\mu E_{h}$ per electron, respectively.
@@ 1 Introduction
Most DBFS generators are designed only for DF in the context of evaluating the Coulomb potential of a charge density. These include the AutoABS method by Yang et al.57, atomic Cholesky decomposition (aCD) and atomic compact Cholesky decomposition (acCD)64, 65, long-range corrected auxiliary basis sets by Hellmann et al.58, even tempered auxiliary basis sets with shared exponents by Daz-Tinoco et al.59 and a similar approach has also recently been reported for relativistic OBSs60.
@@
The AutoAux method, available in the ORCA software package66 and in the command-line interface of Basis Set Exchange (BSE)41, produces a primitive even-tempered DFBS that can be applied universally (in mean-field and correlated simulations). The AutoAux generator uses even-tempered sequences of Gaussian primitive AOs to span the OBS-deduced range of exponents for each angular momentum block. The exponent ranges and even-tempered ratios vary with the angular momenta ($L$) and the atomic number $Z$ (these even-tempering parameters are prescribed for up to $L=7$; the BSE implementation of AutoAux uses the same exponent ratio for $L\geq 7$). The generator also limits the highest angular momentum of DFBS AOs below what is needed for exact atomic DF, in a manner that depends on the highest angular momentum of the occupied orbitals. AutoAux as a result is quite complex, defined with more than a dozen model parameters in total.
@@
PySCF’s simple built-in generator67 produces even-tempered primitive DFBS and is very similar in spirit to AutoAux. It has fewer model parameters (e.g., even-tempering ratio is not graded with angular momenta) and claims to offer similar accuracy to AutoAux, although no extensive comparison exists.
@@
Lehtola presented a DFBS generator that produces nearly exact atomic DFBS composed of primitive solid-harmonic Gaussian (SHG) AOs.62 First, the OBS AOs on a given atom are used to generate a base pool of primitive SHGs by approximating each angular momentum channel of solid-harmonic OBS AO products by a single solid-harmonic AO. The resulting pool is then regularized by pivoted Cholesky decomposition (pCD) of the two-center two-electron Coulomb integral matrix, with the accuracy of the resulting primitive DFBS controlled robustly by a single threshold. Although the resulting DFBSs are nearly exact, they are extremely large, lead to high condition numbers in molecules, and contain primitive Gaussians with twice the angular momentum of the highest angular momentum ($L_{\text{OBS}}$) of functions in the parent OBS. Hellmann et al.68 extended Lehtola’s method62 to include exact spherical functions with mixed angular momentum functions instead of creating multiple angular momentum channels; however, it does not address the large number of primitives in the generated DFBS. Lehtola addressed the latter issue in his recent work69 by extending the algorithm to prune high-angular momentum functions out of the DFBS with a tunable parameter. The algorithm then uses Kállay’s70 contraction method with singular value decomposition (SVD) to contract the DFBS, further reducing its size.
An attractive alternative to density fitting that avoids these issues and provides some of its advantages is the (pivoted) Cholesky decomposition (CD).71 Any positive-definite 2-particle interaction represented in an arbitrary OBS can be decomposed with accuracy controlled by a single threshold. CD has been used for electronic structure simulations in both relativistic72, 73, 74 and nonrelativistic75, 76, 77 contexts and close connections between CD and DF have been demonstrated78. Although the accuracy of CD can be robustly pushed beyond the reach of conventional density fitting, the cost of molecular CD can be substantially higher than that of DF-based approaches and CD is not as universally applicable black-box technology as desired.79 The use of CD for construction of atomic density fitting basis sets, under the name of atomic CD (aCD), was pioneered by Aquilante et al.64 The improved version of aCD, termed atomic compact CD (acCD),80 was proposed later that greatly reduced the number of primitive and contracted DF AOs. Both the aCD and acCD approaches can be viewed as black-box 1-parameter DFBS generators and are also unbiased towards mean-field and correlated methods, albeit they produce deeply contracted DFBS AOs.
@@
To navigate the severe gaps in the coverage by existing manually-constructed DFBS and the notable shortcomings of the existing DFBS generators we attempted to design a new DFBS generator that produces (a) primitive Gaussian DFBSs, (b) of comparable size and accuracy as the manually-optimized DFBS, (c) usable in the context of mean-field and correlated simulation, and (d) has as few adjustable parameters as possible to ensure maximum universality. By combining the existing ideas for spanning the product space of OBS AOs with new ideas for regularizing and pruning the exact DFBS using atomic correlated ensemble RDMs we arrived at a DFBS generator dubbed MADF. The purpose of this manuscript is to describe it and assess its performance. Following a quick recap of density fitting in Section 2.1 we describe the 2 key ingredients of the MADF generator: the regularization of the complete set of candidate primitives (described in Section 2.2) and subsequent pruning by a 2-body energy estimator model (described in Section 2.3). Section 3 describes the technical and details of the parameter training regiment. In Section 4 we document the performance of the DFBSs generated by our method against the manually-optimized DFBSs and the DFBSs generated by the AutoAux algorithm and acCD.
@@ 2.1 Density Fitting
[…]
Such choice makes the error in $(\mu\nu|\hat{O}|\rho\sigma)$ quadratic in the density fitting errors and is a special case of robust density fitting;81, 82 if the fitting norm were defined with operator $\hat{W}\neq\hat{O}$ the explicit robust fitting approach would be necessary and variational extensions83 could be introduced for convenience. Least-squares minimization of $\delta_{\mu\nu}||_{\hat{O}}$ is equivalent to solving the following equation
@@
DFBS generators usually57, 58, 59, 60, 61, 62, 63, 66 work in a single-atom context, i.e., DFBS is constructed using information about products on OBS AOs on a single atom. It is reasonable to question whether this is sufficient. Assuming a system composed of a single atom type, global density fitting in a system of $n$ atoms will be able to fit every such 1-center product of OBS AOs using $n$ times as many DFBS AOs as in a single atom. Indeed, for long-range $\hat{O}$ the fitting of even strongly-localized AO products generally involves DFBS AOs from far away.84 Thus, 1-center products in a molecule can borrow DFBS AOs from other centers to make the fitting more accurate than in a single atom. This suggests that a single atom poses a more challenging DF setting than a molecule, which supports the use of a single-atom setting for the generation of DFBS.
@@ 2.2 Products of primitive Gaussians AOs
The density fitting becomes exact if the span of DFBS matches the span of the OBS AO products. Because DFBS AOs are atom-centered but the products of AOs in a molecule are generally centered between atoms, exact DF is only possible for a single atom. We then confine ourselves to a single atom. Furthermore, we will limit ourselves to the case of OBS and DFBS composed of solid-harmonic Gaussian (SHG) AOs. Unnormalized primitive SHG with exponent $\alpha$ is defined as:85, 86
@@
Whereas product of any number of concentric primitive Cartesian Gaussians is a single Cartesian Gaussian (modulo normalization), product of concentric SHGs $\chi_{\alpha_{\mu},l_{\mu},m_{\mu}}$ and $\chi_{\alpha_{\nu},l_{\nu},m_{\nu}}$ is a linear combination of $r^{l_{\mu}+l_{\nu}-L}\chi_{\alpha_{\mu}+\alpha_{\nu},L,m_{\mu}+m_{\nu}}$ for $\max(|l_{\mu}-l_{\nu}|,m_{\mu}+m_{\nu})\leq L\leq l_{\mu}+l_{\nu}$. Clearly, the contribution to the product from the Clebsch-Gordan channel $L<l_{\mu}+l_{\nu}$ is an SHG multiplied by $r^{l_{\mu}+l_{\nu}-L}$. Following Stoychev61, Lehtola proposed to account for this effect by approximating each such contribution by a single SHG of angular momentum $L$ with the following effective exponent (see Appendix II of Ref. 62):
@@
Let $\mathcal{C}$ denote the “complete” set of all unique solid harmonic primitive AOs generated thereby from all products of primitive OBS AOs on a given atom. The set $\mathcal{C}$ can be used as DFBS for nearly exact density fitting in an atom (a) for any contraction of the generating set of primitive OBS AOs and (b) for any operator. However, even in a single atom such DFBS is highly linearly dependent and overcomplete, 87, 62 therefore, it is unusable with finite precision arithmetic. Moreover, such a basis would be extremely large, defeating the purpose of the DF approximation. Hence, a compact DFBS is needed to represent the product space of AO functions. Since the set $\mathcal{C}$ is overcomplete for use as a DFBS for an atom, the unimportant and linearly dependent functions must be trimmed out of this set. Lehtola proposed to do such pruning by pCD.62 However, even after pCD the resulting DFBSs are still large, as illustrated in Figure 1. While pCD reduces the density of the complete set, it largely leaves its range of exponents intact (at least for lower $L$), and it still keeps many pairs of adjacent exponents with small ratios ($\ll 2$). The problem is especially pronounced for the high angular momentum ($L=6$).
@@
Figure 1: Exponents of DFBS generated (or manually-optimized) for Kr atom with representative nonrelativistic and relativistic OBS using various methods scattered on a logarithmic number line for different angular momenta i.e., $L=\{0,1,2,6\}$. The number of exponents is shown on the right of each scatter plot. Figure 1(a) shows distribution of DFBS exponents for nonrelativistic OBS def2-QZVPP and Figure 1(b) for relativistic OBS x2c-QZVPPall-2c. The plot depict Lehtola’s complete set (shown in purple), Lehtola’s pCD-regularized set (shown in red) obtained using ERKALE88 with pCD threshold $10^{-7}$ as recommended in Ref.62, even-tempered exponents produced by AutoAux (shown in green) and set of exponents produced by the MADF algorithm (shown in orange). Distribution of exponents of manually-optimized DFBS def2-QZVPP-RIFIT is also shown (in blue) in Figure 1(a).
@@
In this work, we regularize the complete candidate exponent pool using a simple adaptive algorithm designed to produce non-even-tempered sets that deviate from the input set as little as possible and with exponent ratios as least as large as the target ratio threshold. Given a primitive SHG set $\mathcal{C}$ and a target exponent ratio threshold $\zeta$, each subset $\mathcal{C}_{L}$ containing primitives of angular momenta $L$ is regularized by repeatedly replacing the pair of primitive SHG shells with the smallest exponent ratio $\alpha_{1}/\alpha_{2}$ by a single primitive SHG shell with the geometric mean exponent $\alpha=\sqrt{\alpha_{1}\alpha_{2}}$. If multiple exponent pairs have a “soft” tie for the smallest exponent ratio, the pair with the smallest orbital exponent wins. Regularization stops if no pairs of primitives have an exponent ratio less than $\zeta$. Clearly, regularization is most easily implemented if $\mathcal{C}_{L}$ is first sorted by the exponent in descending order. The algorithm is described in detail in Algorithm 1. The recommended value of $\zeta$ will be determined heuristically in Section 4.1.
@@
2:  $\mathcal{C}\leftarrow\textsc{CompleteDFBS}(\mathcal{P})$ $\triangleright$ Complete set of DF SHG shells with exponents generated by Eq. 8
3:  $L_{\text{max}}\leftarrow\textsc{MaxAM}(\mathcal{c})$ $\triangleright$ Maximum angular momentum of shells in $\mathcal{c}$
4:  $\mathcal{R}\leftarrow\{\}$ $\triangleright$ Initialize the regularized complete DFBS
@@
6:   $\mathcal{c}\leftarrow\textsc{SelectByAM}(\mathcal{C},L)$ $\triangleright$ Shells in $\mathcal{C}$ with angular momentum $L$
[…]
8:     $\zeta_{\text{min}}\leftarrow\textsc{MaxRatio}(\mathcal{c})$ $\triangleright$ smallest exponent ratio
@@
11:      $\mathcal{c}\leftarrow\textsc{SortByExponent}(\mathcal{c})$ $\triangleright$ Sort in order of decreasing exponents
12:      $\mathcal{p}\leftarrow\textsc{Neighbors}(\mathcal{c})$ $\triangleright$ List of {Exponent($\mathcal{c}[i]$), Exponent($\mathcal{c}[i+1]$), ii} tuples
13:      $\mathcal{p}\leftarrow\textsc{SortByExponentRatio}(\mathcal{p})$ $\triangleright$ Sort in order of increasing exponent ratios
14:      $\zeta_{\text{min}}\leftarrow\mathcal{p}[0,0]/\mathcal{p}[0,1]$
15:      $\mathcal{p}\leftarrow\textsc{MinRatioTies}(\mathcal{p},\zeta_{\text{min}})$ $\triangleright$ Select pairs with exponent ratios in $\zeta_{\text{min}},\zeta_{\text{min}}+10^{-8}]$
16:      $\{\alpha_{I},\alpha_{I+1},I_{\text{min}}\}\leftarrow\textsc{MinElement}(\mathcal{p})$ $\triangleright$ Find the pair with the smallest exponents
17:      $c\leftarrow\textsc{SHGShell}(\sqrt{\alpha_{I}\alpha_{I+1}},L)$ $\triangleright$ Pair of shells with closest exponents fused to single shell
18:      $\mathcal{c}\leftarrow$ Concat($\mathcal{c}\left[0:I_{\text{min}}-1\right],c,\mathcal{c}\left[I_{\text{min}}+2:-1\right]$)
19:     end while
20:   end if
21:   $\mathcal{R}\leftarrow\textsc{Concat}(\mathcal{R},\mathcal{c})$
22:  end for
23:  $\mathcal{R}\leftarrow\textsc{CholeskyScreen}(\mathcal{R},10^{-8})$ $\triangleright$ See Ref.62, only drops strongly linearly dependent primitives.
24:  return $\mathcal{R}$
25:end function